# SlideSLAM: Sparse, Lightweight, Decentralized Metric-Semantic SLAM for Multi-Robot Navigation

### v5

@@ Abstract
[…]
@@
Metric-Semantic SLAM; Multi-Robot Systems; Place Recognition; Aerial Systems; Heterogeneous Robots
@@ I Introduction
Figure 1: Robot platforms used in our experiments. We utilize three types of robots for our experiments. Among these are two aerial platforms: the Falcon 250 UAV (left) and the Falcon 4 UAV (middle), along with one ground platform, the Scarab UGV (right). The Light Detection and Ranging (LiDAR)-equipped robot (Falcon 4) is primarily used for outdoor operations due to its size, and superior sensing capabilities. In contrast, the RGB and Depth (RGBD) camera-based robots (Falcon 250 and Scarab) are more suitable for cluttered indoor environments due to their smaller footprints. All three platforms have GPS-denied autonomous navigation capabilities, enabling them to safely explore cluttered environments using only onboard computation and sensing.
@@
Figure 2: Metric-semantic SLAM results from seven data sequences collected by heterogeneous robots. Trajectories in different colors correspond to different data sequences. 2(a) shows a 3D reconstruction of the Pennovation campus at the University of Pennsylvania and helps visualize the environment. Outdoor objects, such as vehicles, tree trunks, and light poles are mapped as shown in 2(e). Indoor objects, such as chairs, tables, and monitors are mapped as shown in 2(h). 2(b) shows the same metric-semantic map overlayed on top of an accumulated point cloud constructed by our Falcon 4 UAV. 2(c) shows an orthophoto depicting the merged metric-semantic map of three parking lots and two buildings constructed by seven robots. 2(g) and 2(h) show a zoomed-in view of one of the lab buildings and illustrate the metric-semantic qualitative mapping performance in the indoor environments.
@@
Table I: Related work. This table compares SlideSLAM against several state-of-the-art methods, focusing on key attributes essential for enabling semantics-in-the-loop autonomous exploration with robot teams. In this table, decentralized methods support operation with asynchronous observation updates from other robots. Semantic localization means that semantic objects are used for localization between consecutive key poses. Semantic loop closure indicates that semantic objects are explicitly used for place recognition and loop closures. Optimization over object models involves explicitly modeling objects as geometric shapes and optimizing them jointly with robot poses over time. Real-time with autonomy means that the metric-semantic SLAM system operates in real-time on-board the robots and is fully integrated into an autonomous exploration or navigation system. We develop the first metric-semantic SLAM algorithm that supports decentralized operation and is deployed to enable autonomous navigation and exploration using heterogeneous robot teams.
@@
Robotic systems are expected to make an impact in demanding applications, such as forest inventory management, orchard yield estimation, infrastructure inspection and household assistance. This demands interpreting human instructions given in semantically meaningful terms relating to objects and properties in the robot’s environment. To execute such missions autonomously, robots must understand their environment beyond its geometric structure and perceive it at a semantic level. This requires building and maintaining a semantically meaningful representation of the environment that encodes actionable information (e.g., timber volume and health in forests, corrosion in infrastructure, survivors’ locations in natural disasters). Such a representation has to be storage efficient for the robots to maintain over large-scale missions, and has to allow efficient optimization for SLAM.
@@
While traditional SLAM approaches [15, 16] offer excellent accuracy in geometric perception, including estimating robot poses and reconstructing 3D geometric structures (points, surfaces, voxels), they are often insufficient to support large-scale missions, particularly when used with multiple heterogeneous robots for real-time autonomy. The challenges include managing the high computational load associated with long-time-horizon SLAM, the high storage demand for large geometric maps, and handling loop closure and large-scale map merging operations. Additionally, current methods offer limited generalizability across different robot platforms and sensing modalities and lack semantic information in the resulting maps, which is required for executing semantically meaningful tasks.
We develop a multi-robot decentralized metric-semantic SLAM system that is open-sourced [^1] Code: https://github.com/KumarRobotics/SLIDE_SLAM. Project website: https://xurobotics.github.io/slideslam/. and suitable for real-time autonomous navigation and exploration with resource-constrained heterogeneous robots. We model the environment using hierarchical metric-semantic maps containing three levels of abstraction: high-level, mid-level, and low-level. The high-level abstraction stores a sparse landmark map that explicitly models objects and can be used for task planning and long-time-horizon SLAM. The mid-level abstraction maintains an optional accumulated segmented point cloud map that can be activated to provide detailed semantic information on demand. The low-level abstraction is a volumetric map over short durations, which is used for navigation and obstacle avoidance.
@@
This proposed system benefits significantly from the high-level object-level map representation which offers numerous advantages such as: First, it circumvents the impracticality and high computation and memory demands of raw point clouds or images, particularly for large-scale missions with Size, Weight, and Power (SWaP) constrained robot teams. Second, the lightweight nature of our representation ensures that robot-to-robot information sharing is feasible, even with limited communication bandwidth. This, coupled with our semantics-driven place recognition and loop closure algorithm, enables efficient detection of inter-robot loop closures. Third, it provides a direct and intuitive representation for robots to accomplish high-level semantically meaningful tasks, such as actively exploring to search and reduce uncertainties in objects of interest. Lastly, unlike traditional approaches based on dense geometric features that usually have to marginalize variables frequently, which unavoidably leads to loss of information, our representation enables us to keep track of actionable information over a much larger scale.
@@
Algorithm: We develop a real-time decentralized metric-semantic SLAM framework that supports heterogeneous aerial and ground robots. The framework features:
@@
    a flexible front-end that supports LiDAR and RGBD sensors, and can perform both closed-set as well as open-vocabulary object detection,
@@
    an efficient back-end customized for our object-level metric-semantic map representation,
@@
    semantics-driven place recognition algorithms that utilize sparse and lightweight object-level semantic maps for map merging,
  4. (d)
@@
[…]
Figure 3: System Diagram. Our system takes in data streams from each robot’s onboard sensors, which can be either an RGBD camera or a LiDAR, and performs instance segmentation to extract semantic object features. Meanwhile, low-level odometry, either Visual-Inertial Odometry (VIO) [17] or LiDAR-Inertial Odometry (LIO) [18], provides relative-motion estimates between consecutive key poses. Next, the metric-semantic SLAM algorithm takes in such semantic observations and relative motion estimates, and constructs a factor graph consisting of both robot pose nodes and semantic object landmark nodes. Meanwhile, our multi-robot communication module (see Fig. 4) opportunistically leverages communication to share lightweight semantic observations among robots in a decentralized way. Based on this shared information, our metric-semantic place recognition algorithm constantly checks for possible inter-robot loop closures at a fixed rate. Once a loop closure is detected, the resulting transformation between each pair of robots is used to transform all observations into each robot’s reference frame. These observations are then added to their own factor graphs, forming a merged metric-semantic map. Note that the entire perception-action loop runs in a decentralized manner onboard each robot. Besides the obvious differences in control algorithms, the planning modules and the front-end processing algorithms are also different across each robot platform. This is due to the need to accommodate the differences in sensing modalities (RGBD and LiDAR), operating environments (indoor, urban, and forests), and traversal modes (ground and aerial). However, the core metric-semantic SLAM framework remains the same.
@@ II-A Metric-semantic SLAM
Unlike traditional SLAM, metric-semantic SLAM constructs a map that encodes geometric features and semantic information on objects of interest. Metric-semantic SLAM has gained significant success and popularity in the past decade [19], driven by the rapid development of deep learning techniques that extract semantic information from sensor data. A variety of map representations are used in existing metric-semantic SLAM literature. Some use dense semantic maps, such as meshes [20], volumetric maps [21, 22], surfels [23], and 2.5D grid maps [24, 25, 26]. Others use sparse object-level maps with prior information of object shapes, such as centroids [11], cubes [8], ellipsoids [9], cylinders [27], structured object models with prior shape constraints [28, 7], and mesh-based object models [29, 12]. While dense metric-semantic maps are suitable for obstacle avoidance, their computational and memory demands, especially in SWaP-constrained platforms, can be significant. Thus, sparse object-level maps are desirable for real-time downstream tasks such as active information gathering, object manipulation and multi-robot collaboration. This is because semantic information is essential for robots to perceive their environment intelligently and execute tasks accordingly. Sparsity helps robots reduce resource demands on computation, storage, and communication. Therefore, our work seeks to develop a metric-semantic SLAM framework that utilizes sparse explicitly modeled objects and supports real-time decentralized operations within a heterogeneous robot team.
@@ II-B Place recognition
Place recognition and loop closure address the challenge of identifying whether a robot revisits a previously mapped location. This enables odometry drift correction in single-robot systems (intra-robot loop closure) and map merging in multi-robot systems (inter-robot loop closure). Prevalent approaches typically use consistency graphs. These methods operate on the assumption that pairwise distances between points or landmarks remain unchanged across two candidate maps [30, 31, 32, 33]. The largest set of consistent associations is then identified and used to estimate the transformation. However, this assumption is invalid with the object-level map representations, where object detection and localization errors introduce additional inconsistency. Conversely, semantic maps are much sparser and provide richer information. Some approaches leverage this fact and propose descriptor-based methods based on the spatial relationship of semantic objects. In [34], a random walk descriptor encoding semantic labels of neighboring nodes is designed for each semantic node. A semantic histogram-descriptor-based method is proposed in [35], which improves the efficiency of graph matching compared to the random-walk-based method. [4] also leverages random walk descriptors with semantic information but further adds information on object parameters, angles, and distance between an object node and its neighboring nodes to filter out false candidates. In [36], Urquhart tessellations are derived based on the positions of semantic landmarks (tree trunks in forests). Polygon-based descriptors are then computed to describe the local neighborhood of the robot, which are then used for place recognition. Hierarchical descriptors consisting of appearance-based descriptors like Distributed Bag of Words (DBoW2) [37] and higher level descriptors encoding information about a node’s neighboring objects or places are designed in [38, 2]. Drawing inspiration from some of these prior works, we design two complementary place recognition algorithms that leverage our metric-semantic map representation and adopt exhaustive search-based and descriptor-based strategies, respectively.
@@ II-C Multi-robot SLAM
Multi-robot SLAM systems expand upon their single-robot counterparts by incorporating two crucial modules: inter-robot loop closure and multi-agent graph optimization. Numerous studies address the challenge of place recognition, as detailed in II-B. Representative works in multi-robot SLAM include [39, 40], which rely on a central base station or agent to merge measurements from multiple robots, and [41, 42, 43, 44], which operate in a distributed or decentralized manner. Distributed multi-robot systems involve robots that coordinate through message passing, possibly with central nodes, while decentralized systems operate without a central decision point, enhancing robustness by allowing each robot to make independent decisions. Decentralized multi-robot SLAM is a promising framework in multi-robot SLAM research. It enables robots to operate independently and collaborate opportunistically when communication links are established, allowing them to accomplish complex tasks such as large-scale collaborative exploration without relying on infrastructure.
@@ II-D Semantics-in-the-loop navigation and exploration
Figure 5: Semantics-in-the-loop autonomy. The Falcon 250 UAV autonomously explores the environment while constructing the voxel map utilized for low-level navigation and the metric-semantic map used for active semantic loop closure in real time. The long-horizon exploration trajectory (thin green line) is shown alongside the dynamically feasible trajectory (thick blue curve) planned toward the next exploration waypoint.
@@ III Problem Formulation
Given an unknown environment and a team of heterogeneous robots, our objective is to construct a hierarchical metric-semantic map of the environment without relying on any infrastructure. The robots can communicate with each other opportunistically. Each robot must estimate its pose and the state of the environment and use this information to determine the next best actions for navigation and exploration in real time.
@@ III-A Preliminaries
Consider $K$ robots, indexed from 1 through $K$. The map of the $k$-th robot $\mathcal{M}^{(k)}$ consists of a set of objects $\{\mathbf{\ell}^{k}_{1},\ell^{k}_{2},...,\ell^{k}_{n_{k}}\}$ belonging to a set of pre-defined classes (i.e., object categories). An object is described by its class and a state vector, including position, orientation, and shape information. For example, a cuboid has three shape parameters describing the width, length, and height. A cylinder has one shape parameter describing the radius (we do not estimate its length as it is unobservable for the majority of measurements due to limited field of view). Ellipsoid models are simplified to have two shape parameters (radius and height) and known identity orientation. Objects of each class are modeled using one of the aforementioned three shapes as specified by the user. For each key pose of robot $k$, denoted by $\mathbf{x}^{k}_{t}$, the associated measurements include object detections from this pose and the relative motion from the previous key pose. The details of the shape models, robot measurements, and the factor graph used for optimization are presented in Section IV.

### v6

@@ Abstract
This paper develops a real-time decentralized metric-semantic SLAM algorithm that enables a heterogeneous robot team to collaboratively construct object-based metric-semantic maps. The proposed framework integrates a data-driven front-end for instance segmentation from either RGBD cameras or LiDARs and a custom back-end for optimizing robot trajectories and object landmarks in the map. To allow multiple robots to merge their information, we design semantics-driven place recognition algorithms that leverage the informativeness and viewpoint invariance of the object-level metric-semantic map for inter-robot loop closure detection. A communication module is designed to track each robot’s observations and those of other robots whenever communication links are available. The framework supports real-time, decentralized operation onboard the robots and has been integrated with three types of aerial and ground platforms. We validate its effectiveness through experiments in both indoor and outdoor environments, as well as benchmarks on public datasets and comparisons with existing methods. The framework is open-sourced and suitable for both single-agent and multi-robot real-time metric-semantic SLAM applications. The code is available at: https://github.com/KumarRobotics/SLIDE_SLAM.
@@
Metric-Semantic SLAM; Multi-Robot Systems; Aerial Systems: Perception and Autonomy; SLAM
@@ I Introduction
Figure 1: Robot platforms used in our experiments. We utilize three types of robots for our experiments: two aerial platforms, the Falcon 250 UAV (left) and the Falcon 4 UAV (middle), and one ground platform, the Scarab UGV (right). The Light Detection and Ranging (LiDAR)-equipped robot (Falcon 4) is primarily used for outdoor operations due to its size and superior sensing capabilities. The RGB and Depth (RGBD) camera-based robots (Falcon 250 and Scarab) are more suitable for cluttered indoor environments due to their smaller footprints. All three platforms have GPS-denied autonomous navigation capabilities, enabling them to safely explore cluttered environments using only onboard computation and sensing.
@@
Figure 2: Metric-semantic SLAM results from seven data sequences collected by heterogeneous robots. Trajectories in different colors correspond to different data sequences. Fig. 2(a) shows a 3D reconstruction of the Pennovation campus at the University of Pennsylvania. Outdoor objects, such as vehicles, tree trunks, and light poles are mapped as shown in 2(e). Indoor objects, such as chairs, tables, and monitors are mapped as shown in Fig. 2(h). Fig. 2(b) shows the same metric-semantic map overlayed on top of an accumulated point cloud constructed by our Falcon 4 UAV. Fig. 2(c) shows an orthophoto depicting the merged metric-semantic map of three parking lots and two buildings constructed by seven robots. Fig. 2(g) and Fig. 2(h) show a zoomed-in view of one of the lab buildings, and illustrate the metric-semantic qualitative mapping performance in the indoor environments.
Table I: Related work. This table compares SlideSLAM against several state-of-the-art methods, focusing on key attributes for enabling semantics-in-the-loop autonomous exploration with robot teams. In this table, decentralized methods support operation with asynchronous observation updates from other robots. Semantic localization means that objects are used for localization between consecutive key poses. Semantic loop closure indicates that objects are explicitly used for place recognition and loop closures. Optimization over object models involves explicitly modeling objects as geometric shapes and optimizing them jointly with robot poses over time. Real-time with autonomy means that the metric-semantic SLAM system operates in real-time on-board the robots and is fully integrated into an autonomous exploration or navigation system. We develop the first metric-semantic SLAM algorithm that supports decentralized operation and is deployed to enable autonomous navigation and exploration using heterogeneous robot teams.
@@
Robotic systems are expected to make an impact in demanding applications, such as forest inventory management, orchard yield estimation, infrastructure inspection and household assistance. This demands interpreting human instructions given in semantically meaningful terms relating to objects and properties in the robot’s environment. To execute such missions autonomously, robots must understand their environment beyond its geometric structure and perceive it at a semantic level. This requires building and maintaining a semantically meaningful representation of the environment that encodes actionable information (e.g., timber volume and health in forests, corrosion in infrastructure, survivors’ locations in natural disasters). Such a representation has to be storage efficient for the robots to maintain over large-scale missions, and has to allow efficient optimization for Simultaneous Localization and Mapping (SLAM).
@@
While traditional SLAM approaches [15, 16] offer excellent accuracy in geometric perception, including estimating robot poses and reconstructing 3D geometric structures (points, surfaces, voxels), they are often insufficient to support large-scale missions, particularly when used with multiple heterogeneous robots for real-time autonomy. The challenges include managing the excessive computational load, the high storage demand for large-scale maps, handling loop closure and map merging operations. Moreover, they offer limited generalizability across different robot platforms or sensing modalities and lack semantically meaningful map representations, which are required for executing semantically meaningful tasks.
Motivated by this, recent advances in SLAM have pushed toward constructing metric-semantic maps [11, 9, 8, 17, 18]. However, they are typically limited to single-robot setups. For scalability and efficiency, an ideal system should support deployment across heterogeneous robot teams, enabling effective collaboration without relying on centralized communication or global localization infrastructure. Extending single-robot metric-semantic SLAM to decentralized operation across multiple heterogeneous robots introduces several unique challenges. These include efficient place recognition and map merging among robots, maintaining consistent and probabilistically sound sensor measurement fusion under intermittent communication, and enabling collaboration among heterogeneous platforms.
@@
To address these challenges, we develop a multi-robot decentralized metric-semantic SLAM system that is open-sourced [^1] Code: https://github.com/KumarRobotics/SLIDE_SLAM. Project website: https://xurobotics.github.io/slideslam/. and suitable for real-time autonomous navigation and exploration onboard resource-constrained single-robot as well as heterogeneous multi-robot platforms. We enable this by mapping the environment using a sparse map representation that explicitly models objects using simple shapes and can be used for task planning and long-horizon SLAM. The proposed sparse object-level map representation offers numerous advantages. First, it avoids the high runtime and memory demands of using raw point clouds or images, which is particularly important for large-scale missions with Size, Weight, and Power (SWaP) constrained robot teams. Second, the lightweight nature of our representation ensures that robot-to-robot information sharing is feasible, even with limited communication bandwidth. This, coupled with our semantics-driven place recognition and loop closure algorithm, enables efficient detection of inter-robot loop closures. Third, it provides a direct and intuitive representation for robots to accomplish high-level semantically meaningful tasks, such as actively exploring to search and reduce uncertainties in objects of interest. Fourth, unlike traditional approaches based on dense geometric features that usually have to marginalize variables frequently, which unavoidably leads to loss of information, our representation enables us to keep track of actionable information over a much larger scale. Finally, it provides a unified representation across different sensing modalities.
@@
Algorithm: We develop a real-time decentralized metric-semantic SLAM framework that supports heterogeneous aerial and ground robots, which includes
@@
    a computationally efficient back-end that uses our object-level metric-semantic map representation, and a flexible front-end that supports both LiDAR and RGBD sensors,
@@
    semantics-driven place recognition algorithms that use sparse object-level maps for map merging,
@@
Experiments: We conduct extensive real-world experiments and provide thorough empirical results and analysis that highlight the efficiency, accuracy, and robustness of our system. We have also made our framework available to the public.
Figure 3: System Diagram. Our system takes in data streams from each robot’s onboard sensors, which can be either an RGBD camera or a LiDAR, and performs instance segmentation to extract semantic object features. Meanwhile, low-level odometry, either Visual-Inertial Odometry (VIO) [19] or LiDAR-Inertial Odometry (LIO) [20], provides relative-motion estimates between consecutive key poses. Next, the metric-semantic SLAM algorithm takes in such semantic observations and relative motion estimates, and constructs a factor graph consisting of both robot pose nodes and object landmark nodes. Meanwhile, our multi-robot communication module (see Fig. 4) opportunistically leverages connectivity to share lightweight semantic observations among robots in a decentralized way. Based on this shared information, our metric-semantic place recognition algorithm constantly checks for possible inter-robot loop closures at a fixed rate. Once a loop closure is detected, the resulting transformation between each pair of robots is used to transform all observations into each robot’s reference frame. These observations are then added to their own factor graphs, forming a merged metric-semantic map. Note that the entire perception-action loop runs in a decentralized manner onboard each robot. Besides the obvious differences in control algorithms, the planning modules and the front-end processing algorithms are also different across each robot platform. This is due to the need to accommodate the differences in sensing modalities (RGBD and LiDAR), operating environments (indoor, urban, and forest), and traversal modes (ground and aerial). However, the core metric-semantic SLAM framework remains the same.
@@ II-A Metric-semantic SLAM
Unlike traditional SLAM, metric-semantic SLAM constructs a map that encodes geometric features and semantic information on objects of interest. Metric-semantic SLAM has gained significant success and popularity in the past decade [21], driven by the rapid development of deep learning techniques that extract semantic information from sensor data. A variety of map representations are used in existing metric-semantic SLAM literature. Some use dense semantic maps, such as meshes [17], volumetric maps [22, 23], surfels [24], and 2.5D grid maps [25, 26, 27]. Others use sparse object-level maps with prior information of object shapes, such as centroids [11], cubes [8], ellipsoids [9], cylinders [28], structured object models with prior shape constraints [29, 7], and mesh-based object models [30, 12]. Following the recent trends in neural implicit representation, some prior works [31, 32, 33] have also incorporated it into a map representation capable for metric-semantic SLAM, where objects are modeled implicitly using latent features from neural networks and used as localization constraints and object association. While dense metric-semantic maps are suitable for obstacle avoidance and facilitate more fine-grained modeling of objects, their computational and memory demands, especially in SWaP-constrained platforms, can be significant. Thus, sparse object-level maps are desirable for real-time downstream tasks such as active information gathering, object manipulation and multi-robot collaboration. This is because semantic information is necessary for robots to perceive their environment and execute tasks in a semantically meaningful manner. Sparsity helps robots reduce resource demands on computation, storage, and communication. A recent work in this space is ConceptGraphs [18], which builds open-vocabulary, semantically rich 3D object maps. However, ConceptGraphs is not designed for real-time, onboard operations on resource-constrained robots. Its detection and 3D object mapping pipelines require extensive computational resources. By contrast, our work seeks to develop a metric-semantic SLAM framework that utilizes sparse explicitly modeled objects and supports real-time decentralized operations within a heterogeneous robot team.
@@ II-B Place recognition
Place recognition and loop closure address the challenge of identifying whether a robot revisits a previously mapped location. This enables odometry drift correction in single-robot systems (intra-robot loop closure) and map merging in multi-robot systems (inter-robot loop closure). Prevalent approaches typically use consistency graphs. These methods operate on the assumption that pairwise distances between points or landmarks remain unchanged across two candidate maps [34, 35, 36, 37]. The largest set of consistent associations is then identified and used to estimate the transformation. However, this assumption is invalid with the object-level map representations, where object detection and localization errors introduce additional inconsistency. Some geometric-based place recognition methods, such as [38], design hand-crafted features for LiDAR point clouds that are translation and rotation invariant to match them accurately and establish robust place recognition. Other methods like [39] generate similar global matching features and perform relative pose refinement using deep neural networks. While these methods demonstrate impressive place recognition performance, they may have high memory usage when storing a large number of geometric features or require large amounts of data to fine-tune the neural network for place recognition in novel environments. Conversely, semantic maps are much sparser, provide richer information and can be abstracted from any sensing modality, LiDAR or camera. Some approaches leverage this fact and propose descriptor-based methods based on the spatial relationship of semantic objects. In [40], a random walk descriptor encoding semantic labels of neighboring nodes is designed for each semantic node. A semantic histogram-descriptor-based method is proposed in [41], which improves the efficiency of graph matching compared to the random-walk-based method. [4] also leverages random walk descriptors with semantic information but further adds information on object parameters, angles, and distance between an object node and its neighboring nodes to filter out false candidates. In [42], Urquhart tessellations are derived based on the positions of semantic landmarks (tree trunks in forests). Polygon-based descriptors are then computed to describe the local neighborhood of the robot, which are then used for place recognition. Hierarchical descriptors consisting of appearance-based descriptors like Distributed Bag of Words (DBoW2) [43] and higher level descriptors encoding information about a node’s neighboring objects or places are designed in [44, 2]. Drawing inspiration from some of these prior works, we design two complementary place recognition algorithms that leverage our metric-semantic map representation from diverse sensing modalities (LiDAR and RGBD Camera) and adopt exhaustive search-based and descriptor-based strategies, respectively.
@@ II-C Multi-robot SLAM
Multi-robot SLAM systems expand upon their single-robot counterparts by incorporating two crucial modules: inter-robot loop closure and multi-agent graph optimization. Numerous studies address the challenge of place recognition, as detailed in Section II-B. Representative works in multi-robot SLAM include [45, 46], which rely on a central base station or agent to merge measurements from multiple robots, and [47, 48, 49, 50], which operate in a distributed or decentralized manner. Distributed multi-robot systems involve robots that coordinate through message passing, possibly with central nodes, while decentralized systems operate without a central decision point, enhancing robustness by allowing each robot to make independent decisions. Decentralized multi-robot SLAM is a promising framework in multi-robot SLAM research. It enables robots to operate independently and collaborate opportunistically when communication links are established, allowing them to accomplish complex tasks such as large-scale collaborative exploration without relying on infrastructure.
@@ III Problem Formulation
Given an unknown environment and a team of heterogeneous robots, our objective is to construct a hierarchical metric-semantic map of the environment without relying on any infrastructure. The robots can communicate with each other opportunistically. Each robot must estimate its pose and the state of the environment and use this information to determine actions for navigation and exploration in real time.
@@ III-A Preliminaries
Consider $K$ robots, indexed from 1 through $K$. The map of the $k$-th robot $\mathcal{M}^{(k)}$ consists of a set of object landmarks $\{\mathbf{\ell}^{k}_{1},\ell^{k}_{2},\ldots,\ell^{k}_{n_{k}}\}$ belonging to a set of pre-defined classes (i.e., object categories). An object is described by its class and a state vector, including its position, orientation, and shape information. For example, a cuboid has three shape parameters describing its width, length, and height. A cylinder has one parameter describing its radius (we do not estimate its height as it is unobservable for the majority of measurements due to limited field of view). Ellipsoid models are simplified to have two shape parameters (radius and height) and known identity orientation. Objects of each class are modeled using one of the aforementioned three shapes as specified by the user. For each key pose of robot $k$, denoted by $\mathbf{x}^{k}_{t}$, the associated measurements include object detections from this pose and the relative motion from the previous key pose. The details of the shape models, robot measurements, and the factor graph used for optimization are presented in Section IV.